L^{A}TEX News

Issue 8, December 1997

New supported font encodings

Two new font encodings are supported as options to the fontenc package:

OT4 This is a seven-bit encoding designed for Polish. The L^{A}TEX support was developed by Mariusz Olko. TS1 This is the ‘Text Companion Encoding’; it contains symbols designed to be used in text, as opposed to mathematical formulas, and some accents designed for uppercase letters. It is currently supported by the ‘tc’ fonts, which match the T1 encoded ‘ec’ text fonts. A subset of the glyphs in this encoding is supported by virtual fonts distributed with the PostScript font metrics on the ctan archives. (This is the ‘8c’ encoding in Karl Berry’s fontname scheme.) The textcomp package provides access to this encoding but here is a warning to current users of that package: some of the internal names for the characters have changed.

New input encodings

These additions to the inputenc package are decmulti (the DEC Multinational Character Set, contributed by M. Y. Chartoire) and cp1250 (an MS-Windows encoding for Central and Eastern Europe, contributed by Marcin Woliński). There is also a cp1252 encoding that is identical to ansinew.

Tools

The calc package (used in many examples in The L^{A}TEX Companion) has been contributed to this distribution by Kresten Krab Thorup and Frank Jensen. This is essentially the same as the version that has been available from the ctan archives for some time, with one minor change: to use L^{A}TEX-style error messages. It enables the use of arithmetic expressions within arguments to standard L^{A}TEX commands where a length or a counter value is required. For example:

  \setcounter {page} { \value{page} * 2 + 1 }
  \parbox { 3in - ( 2mm + \textwidth / 9 ) }

There have also been some improvements to several other packages in this collection. In particular, bm now works correctly with constructions such as \bm{f'} involving ' or other characters which use TEX’s special “\mathcode"8000” feature. Also, multicol sets the length \columnwidth to an appropriate value; this enables it to work with classes that support two-column setting, e.g., the AMS classes.

Graphics

The special oztex.def driver file has been removed, and OzTEX support has been merged with dvips, following advice from Andrew Trevorrow about OzTEX 3.x. The keyval package has had some internal

improvements: to use L^{A}TEX format error messages; and to avoid ‘# doubling’. This latter change means that the command key for the graphicx version of \includegraphics should now be used with one # rather than two. For example, command = `gunzip #1. Fortunately this key is almost never used in practice, so few if any documents should be affected by this change.

L^{A}TEX3 experimental programming conventions

As announced at the TEX Users Group meeting (Summer 1997), a group of highly experimental packages will soon be released to allow experienced TEX programmers to experiment with, and comment on, a proposed set of syntax conventions and basic data-types that might form the basis for programming large scale projects in TEX. They will be located in this CTAN directory:

                                                     CTAN:macros/latex/packages/expl3

The documentation of this material is as follows: individual package files provide outline, draft documentation; there is an article that gives an overview of the syntax and related concepts; there is a readme.txt file containing a brief description of the collection.

All aspects of these packages are liable, indeed likely, to change. They should not be used at this stage for anything that requires a stable system. However, we do encourage people to experiment with these packages, and to send comments on them to the LaTeX-L mailing list. To subscribe to this list, mail to:

                                                     listserv@urz.uni-heidelberg.de

the following one line message:

                                                     subscribe LATEX-L <first-name> <second-name>

L^{A}TEX News, and the L^{A}TEX software, are brought to you by the L^{A}TEX3 Project Team; Copyright 1997, all rights reserved.